OpenStudy (anonymous):

I need a function table for y=(7/2)^x

OpenStudy (anonymous):

\[y=\left( \frac{ 7 }{ 2 } \right) ^{2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

ty very much

OpenStudy (anonymous):

but when i go into negative 1 etc. how would it start?

OpenStudy (anonymous):

i just need -1

OpenStudy (anonymous):

negative would be the reciprocal 1/(7/2)^x (plug in positive values of x)

OpenStudy (anonymous):

so for x=-1 it would be 2/7

OpenStudy (anonymous):

2/7 and 4/49?

OpenStudy (anonymous):

for -1 and -2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

or you can just plug in negative exponent into your calculator

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

then to get a correct fnction table for a equation like y= (5/4) if x = then y= 5/4 if x = 2 then y = 10/8 and x=-2 then y=8/10 and half of that if x=-1

OpenStudy (anonymous):

can you retype the equation is it y = 5/4 x?

OpenStudy (anonymous):

yes y=(5/4)^x

OpenStudy (anonymous):

sorry i hate this laptop

OpenStudy (anonymous):

it would be the same procedure as in the first example

OpenStudy (anonymous):

so for x=1 y=5/4, x=-1 y=4/5, x=0 y=1

OpenStudy (anonymous):

ok thank you so much have a huge algebra test today actually its 2 am holy ****

OpenStudy (anonymous):

well good luck, get some sleep

OpenStudy (anonymous):

for sure thanks for the help and the luck ima need it lol